The statement “When an integer is added to itself, the sum is greater than the integer” is (A) always true (B) never true (C) true only when the integer is positive (D) true for non-negative integers

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (C) is true i.e only when the integer is positive
For example : consider the positive integer 5 = 5 + 5 = 10
In positive integer the sum is greater than the integer.
But in negative integer -4 = -4 + (-4)

=-4–4

=-8
In negative integer the sum is less than the integer.

Thanks